# Off-Road Navigation via Implicit Neural Representation of Terrain Traversability

### v2

@@ Abstract
Autonomous off-road navigation requires robots to estimate terrain traversability from onboard sensors and plan accordingly. Conventional approaches typically rely on sampling-based planners such as MPPI to generate short-term control actions that aim to minimize traversal time and risk measures derived from the traversability estimates. These planners can react quickly but optimize only over a short look-ahead window, limiting their ability to reason about the full path geometry, which is important for navigating in challenging off-road environments. Moreover, they lack the ability to adjust speed based on the terrain bumpiness, which is important for smooth navigation on challenging terrains. In this paper, we introduce TRAIL (Traversability with an Implicit Learned Representation), an off-road navigation framework that leverages an implicit neural representation to continuously parameterize terrain properties. This representation yields spatial gradients that enable integration with a novel gradient-based trajectory optimization method that adapts the path geometry and speed profile based on terrain traversability.
@@ I Introduction
Autonomous navigation in off-road environments requires capabilities beyond obstacle avoidance. Unlike urban roads, off-road terrain exhibits highly variable and unstructured conditions such as uneven surfaces and vegetation. Consequently, there has been growing interest in developing autonomous systems capable of operating reliably in these settings [fan2021step, jian2022putn, xue2023traversability, dixit2024step, frey2024roadrunner, patel2024roadrunner, yoo2024traversability, lee2025trg, frey2023fast, cai2023probabilistic, cai2024evora, cai2025pietra, mattamala2025wild, gasparino2022wayfast, gasparino2024wayfaster].
Most existing off-road navigation frameworks estimate terrain traversability from either terrain geometry [fan2021step, dixit2024step, frey2024roadrunner, patel2024roadrunner, yoo2024traversability, jian2022putn, xue2023traversability] or vehicle traction on the terrain [frey2023fast, gasparino2024wayfaster, cai2023probabilistic, mattamala2025wild, gasparino2022wayfast, cai2024evora, cai2025pietra], from which a cost map is derived and represented as a grid map. Methods that rely purely on geometric information can be too conservative as they may struggle to distinguish between drivable vegetation, such as tall grass, and rigid obstacles, such as rocks, since both may appear as regions with high elevation and steep slopes. On the other hand, methods that only rely on traction estimate suffer from generalizability, since the supervision signal can only be generated along the vehicle’s traversed path. Moreover, data must be collected in non-traversable (i.e., possibly dangerous) regions – for example, a robot must actually collide with trees enough times to learn that trees are obstacles. On the planning side, gradient-free sampling-based trajectory optimization techniques such as MPPI [williams2017information] are applied to generate a path using the cost grid map. While MPPI offers great flexibility in defining the objective function, it often produces non-smooth, oscillatory control signals [andrejev2025pi], struggles with handling hard constraints [park2025csc], and can require extensive tuning of sampling parameters and weights [homburger2025optimality]. Moreover, these formulations do not reason about the full path geometry or speed profile along the trajectory, which are crucial for smooth navigation in challenging off-road environments. Failing to consider these longer-term factors can lead to unnecessary hardware stress and increased odometry drift, both of which negatively impact downstream planning and control.
@@
In this work, we address limitations in both perception and planning to achieve more reliable off-road navigation. We estimate both geometric properties and terrain-response signals, and incorporate these complementary sources of information into planning and trajectory optimization. Geometric properties of the terrain provide a conservative yet physically grounded prior, while the terrain-response signals capture how the vehicle actually interacts with the terrain. Our method leverages implicit neural representations [peng2020convolutional] to unify perception and planning within a gradient-based trajectory optimization framework, enabling the planner to jointly adapt the path geometry and speed profile to challenging terrain while respecting constraints. In summary, the contributions of our work are:
  • A learning-based implicit neural representation that continuously parametrizes both terrain geometric properties and terrain-response signals. This implicit representation provides smooth spatial structure and readily accessible gradients through a lightweight decoder.
  • A novel gradient-based trajectory optimization framework that directly exploits gradient information from the implicit representation to jointly adapt the path geometry and speed profile according to terrain traversability, while naturally handling constraints.
  • Comprehensive investigations in both simulation and challenging real-world off-road navigation scenarios, demonstrating the effectiveness of the proposed approach compared to conventional frameworks.
II Proposed Method
Given an RGB image and a LiDAR point cloud from onboard sensors, the task is to infer continuous terrain attributes over a local region $x_{\min},x_{\max}]\times[y_{\min},y_{\max}]$ in the vehicle frame and use them to navigate to goals provided by some higher-level planner (which we assume to be given).
II-A Learning Implicit Representations
We train a DNN that maps the image and point cloud to traversability-relevant terrain properties. The network consists of a shared encoder that fuses the inputs into a bird’s-eye-view (BEV) feature plane, followed by multiple decoders. Each decoder takes a query point and its corresponding BEV feature and outputs a continuous-valued terrain property (e.g., maximum ground elevation). Unlike prior work (e.g. [gasparino2024wayfaster, cai2024evora, frey2024roadrunner]), which predicts fixed-resolution grid maps, our implicit formulation yields lightweight decoders, faster inference (Fig. 4), and direct access to spatial gradients for trajectory optimization.
II-A1 Encoder
Our encoder combines Lift-Splat-Shoot (LSS) [philion2020lift] and PointPillars [lang2019pointpillars]. Given an image $X\in\mathbb{R}^{C\times H\times W}$, we define d as the desired downsampling factor. We extract visual features using DINOv2 [oquab2023dinov2], obtaining $X_{\texttt{dino}}\in\mathbb{R}^{C_{\texttt{dino}}\times H_{p}\times W_{p}}$. These features are compressed with a lightweight 2D convolution block ($\operatorname{Conv}+\operatorname{GroupNorm}+\operatorname{ReLU}$) to produce $X_{\texttt{dc}}\in\mathbb{R}^{C_{\texttt{dc}}\times H_{p}\times W_{p}}$. From $X_{\texttt{dc}}$, we predict depth logits $X_{\texttt{depth}}\in\mathbb{R}^{D\times H_{d}\times W_{d}}$, where $H_{d}=H/\texttt{d}$ and $W_{d}=W/\texttt{d}$. We also interpolate $X_{\texttt{dc}}$ to obtain 2D image features $X_{\texttt{feat}}\in\mathbb{R}^{C_{\texttt{dc}}\times H_{d}\times W_{d}}$. The pair $(X_{\texttt{feat}},X_{\texttt{depth}})$ is lifted to a 3D volumetric feature of size $C_{\texttt{dc}}\times D\times H_{d}\times W_{d}$ via depth unprojection, which is then splatted onto the BEV plane, following LSS. The resulting image BEV features are compressed to $X_{\texttt{img}}\in\mathbb{R}^{C_{\texttt{img}}\times Y\times X}$. For LiDAR, we apply PointPillars to obtain point cloud BEV features $X_{\texttt{pcd}}\in\mathbb{R}^{C_{\texttt{pcd}}\times Y\times X}$. The image and point cloud BEV features are concatenated and passed through a 2D convolution block to produce $C_{\texttt{bev}}$ channels, followed by a CBAM attention module [woo2018cbam]. The final encoder output is a fused BEV feature map $X_{\texttt{bev}}\in\mathbb{R}^{C_{\texttt{bev}}\times Y\times X}$.
@@ II-A1 Encoder
Figure 2: Our perception module takes a camera image and a LiDAR point cloud and produces an implicit representation of various terrain properties (only the bump decoder is shown). The proposed planning and control framework then incorporates gradient information provided by the implicit representation into a novel gradient-based trajectory optimization framework, where speed profile and path geometry are co-optimized.
II-A2 Decoder
Denote the input to the encoder as $\mathbf{X}$ (i.e. an image and a point cloud). Given a query point $p=(p_{x},p_{y})$, we can obtain the BEV feature $\psi(p,\mathbf{X})\in\mathbb{R}^{C_{\texttt{bev}}}$ at $p$ from $X_{\texttt{bev}}$. Note that $\psi(p,\mathbf{X})$ is differentiable with respect to $p$, and that $\psi(p,\mathbf{X})$ contains information about neighboring locations as a result of spatial context aggregation by the 2D convolution blocks and CBAM. Then, inspired by [peng2020convolutional], we condition the decoder network on both the query point $p$ and the feature $\psi(p,\mathbf{X})$, thus obtaining an implicit representation (also referred to as a coordinate-based representation). However, decoders that take $p$ directly as input can fail to predict high-frequency features [tancik2020fourier] such as small rocks or partially buried tree roots, which are important for traversability estimation tasks. Instead, we first obtain Fourier features [tancik2020fourier] from $p$, denoted $\gamma(p)\in\mathbb{R}^{\texttt{ff\_dim}}$, and then pass $\gamma(p)$ to the decoder, where ff_dim denotes the dimension of the Fourier feature. This approach has been shown to improve the performance of coordinate-based MLPs on capturing high-frequency features [tancik2020fourier]. To simplify notation, we integrate the BEV feature sampling, Fourier feature mapping, and MLP decoder into a single decoder function $f_{\theta}:\mathbb{R}^{2}\to\mathbb{R}$, defined over the query point $p$. In this formulation, the query point is first used to extract the corresponding BEV and Fourier features, which are then jointly processed by the MLP decoder, i.e. $p\mapsto(\gamma(p),\psi(p,\mathbf{X}))\mapsto\operatorname{MLP}(\gamma(p),\psi(p,\mathbf{X}))$. Our architecture includes five decoders: $f_{\texttt{g\_min}},f_{\texttt{g\_max}},f_{\texttt{c\_mask}},f_{\texttt{c\_min}},f_{\texttt{bump}}$. Two of them estimate the minimum and maximum ground elevations, while two more predict a ceiling mask—indicating the presence of overhanging structures—and the minimum ceiling height, inspired by [meng2023terrainnet]. The ceiling-related decoders are used exclusively during training to prevent ceilings from being misidentified as ground, thereby avoiding inflated ground elevation estimates. The fifth decoder predicts terrain bumpiness, as defined in Section II-A3.
II-A3 Data Preparation and Training
The ground truth labels for ground elevations and ceiling predictions are generated in similar fashion to TerrainNet [meng2023terrainnet], where LiDAR scans in the time range $t-\delta t,t+\delta t]$ are aggregated using odometry to generate dense annotations at time $t$. Here we use $\delta t=15$\mathrm{s}$\$. To quantify terrain bumpiness, we use the root-mean-square (RMS) of the vehicle’s vertical acceleration. A qualitative comparison between this measure (with a window length of $0.1$\mathrm{s}$\$) and traction (defined as the ratio between estimated and commanded velocity [gasparino2022wayfast]) is shown in Figure 3. Note that the RMS of vertical acceleration on the same terrain increases approximately proportionally to the vehicle speed [pereira2009robot]. During dataset generation, we normalize the measured RMS of vertical acceleration by the vehicle speed to compensate for this effect, which is then rescaled to $(0,1)$ using $\operatorname{sigmoid}$. We call the resulting value bumpiness. In our trajectory optimization framework (see Section II-B2), this proportional relationship is intentionally exploited to encourage the optimizer to adapt the speed profile based on predicted terrain bumpiness. To balance the bumpiness data distribution, we compute a histogram of the bumpiness values and give each bumpiness label a weight that is inversely proportional to its frequency. Finally, for training, instead of using another decoder to predict the terrain slope, we directly leverage the differentiability of $f_{\texttt{g\_max}}$. We use PyTorch’s autograd functionality to compute the gradient of $f_{\texttt{g\_max}}$ with respect to the query point, $\partial_{p}f_{\texttt{g\_max}}$, and compute the loss by comparing $\partial_{p}f_{\texttt{g\_max}}$ to the ground truth terrain slope. $L_{1}$ loss is used for $f_{\texttt{g\_min}},f_{\texttt{g\_max}},f_{\texttt{c\_min}},f_{\texttt{bump}}$ and $\partial_{p}f_{\texttt{g\_max}}$. Cross entropy loss is used for both ceiling mask and depth logits prediction.
Figure 3: Comparison of vertical acceleration from an IMU and low-traction events. Peaks in vertical acceleration generally coincide with periods of reduced traction. However, occasional false positives occur due to noisy velocity estimates, leading to misclassified high-traction regions (see right). Camera images are tone- and color-corrected for visibility.
II-B Planning with Implicit Representations
Our planning stack couples a path planner with a trajectory optimizer that exploits spatial gradients from $f_{\texttt{bump}}$. The path planner first generates a coarse path based on geometric properties of the terrain. This path is then refined by the trajectory optimizer, which leverages terrain bumpiness to adjust both the path geometry and the speed profile, reducing unnecessary detours and promoting smoother, terrain-aware motion. This stands in contrast to recent off-road navigation frameworks [gasparino2022wayfast, gasparino2024wayfaster, cai2023probabilistic, cai2024evora, cai2025pietra, frey2024roadrunner], which do not reason about full-path structure and speed adaptation.
II-B1 Path Planner

### v3

# arXiv:2511.18183v3[cs.RO] 27 Feb 2026

@@ Abstract
Autonomous off-road navigation requires robots to estimate terrain traversability from onboard sensors and plan motion accordingly. Conventional approaches typically rely on sampling-based planners such as MPPI to generate short-term control actions that aim to minimize traversal time and risk measures derived from the traversability estimates. These planners can react quickly but optimize only over a short look-ahead window, limiting their ability to reason about the full path geometry, which is important for navigating in challenging off-road environments. Moreover, they lack the ability to adjust speed based on the terrain-induced vibrations, which is important for smooth navigation on challenging terrains. In this paper, we introduce TRAIL (Traversability with an Implicit Learned Representation), an off-road navigation framework that leverages an implicit neural representation to model terrain properties as a continuous field that can be queried at arbitrary locations. This representation yields spatial gradients that enable integration with a novel gradient-based trajectory optimization method that adapts the path geometry and speed profile based on terrain traversability.
@@ I Introduction
Autonomous navigation in off-road environments requires capabilities beyond obstacle avoidance. Unlike structured roads, off-road terrain is unstructured and highly variable (e.g., uneven surfaces and vegetation), motivating substantial recent progress in traversability-aware perception and planning [11, 21, 50, 10, 15, 35, 51, 26, 14, 4, 3, 5, 30, 17, 16, 31, 18, 40, 32, 7].
Most existing off-road navigation stacks derive a grid-based traversability cost map from either exteroceptive terrain cues (e.g., geometry or appearance) or proprioceptive interaction signals (e.g., slip or vibration). Each cue alone has limitations: geometry or appearance-only estimates can be ambiguous in unstructured terrain (e.g., drivable vegetation versus rigid obstacles), while interaction-based supervision is inherently sparse because it is only observed along the vehicle’s traversed trajectories. On the planning side, many systems execute short-horizon, sampling-based control (e.g., MPPI [47]) over these costs; while reactive and flexible, such approaches have limited ability to refine full-path geometry and to adapt a speed profile to terrain-induced disturbances.
In this work, we address limitations in both perception and planning by combining complementary terrain cues. We estimate (i) geometric properties that provide a conservative yet physically grounded prior and (ii) terrain-response signals that reflect the vehicle’s experienced disturbance. We represent these quantities with an implicit neural representation (INR)—a continuous field that can be queried at arbitrary locations—which provides spatial gradients enabling a gradient-based trajectory optimization method that jointly refines path geometry and speed profile subject to various constraints. In summary, our contributions are:
  • A learning-based implicit neural representation that models terrain geometry and terrain-response signals as continuous, queryable functions, with readily accessible spatial gradients through lightweight decoders.
  • A gradient-based trajectory optimization framework that exploits these gradients to co-optimize path geometry and speed profile according to terrain traversability, while naturally handling constraints.
  • Comprehensive evaluations in simulation and challenging real-world off-road experiments demonstrating improved navigation performance over conventional frameworks.
@@
II Related Work
II-A Traversability Estimation
Most existing off-road navigation frameworks estimate terrain traversability based on one of the three approaches:
1) Geometry-based and heuristic methods [11, 10, 15, 35, 51, 21, 50]: These methods estimate traversability from geometric/heuristic cues (e.g., slope, roughness, curvature) and sometimes observability cues (e.g., point sparsity)—often derived from LiDAR—either via handcrafted costs or learned predictors trained with stack-generated pseudo-labels (e.g., RoadRunner [15, 35]). For example, PUTN [21] computes terrain slope, flatness, and point cloud sparsity and fuses them into a single cost for downstream planning. Many works in this category rely mainly on geometric information of the terrain, which may result in conservative traversability estimates (e.g. tall grass may receive a similar cost to rocks).
2) Appearance-based methods [31, 18, 40, 32]: These methods infer traversability from semantic or appearance cues via terrain classification or segmentation, and then the predicted classes can be mapped to planning costs. For example, Shaban et al. [40] train a neural network to classify LiDAR point clouds into discrete traversability classes, which are subsequently converted into a cost map for planning. Works in this category can require substantial manual labeling and may be limited by discrete class definitions.
3) Interaction-based methods [14, 16, 4, 30, 3, 5, 17, 7]: These methods estimate terrain traversability using supervision derived from vehicle–terrain interaction signals (e.g., IMU vibration, slip, suspension). For example, WayFASTER [16] defines traction as the ratio of achieved speed to commanded speed and trains a neural network to map sensor observations to a grid map of traction values. However, a limitation of these methods is that labels are only available along the vehicle’s traversed trajectories, leading to sparse supervision signals and potentially limited generalization. To improve this, recent works [14, 43, 30, 41] leverage vision foundation models (VFMs), such as DINOv2 [34], to obtain general-purpose visual features first and learn mappings from these features (or some compressed version of these features [41]) to traversability estimates.
Relation to this work: In this work, we learn an INR that jointly models terrain geometry and terrain-response signals. The resulting continuous, queryable field provides spatial gradients that our planner exploits for trajectory optimization.
II-B Planning
UGV planning has been extensively studied in both indoor and outdoor environments [12, 23, 24, 6, 49, 27]. Common approaches include graph- and sampling-based planners (e.g., A/D, RRT, and state lattices) [23, 21, 49, 27, 12, 20], as well as trajectory optimization with differentiable map representations (e.g., neural/implicit fields) [24, 6]. In more challenging unstructured settings, planning is often coupled with risk-aware objectives; STEP [11, 10] formulates planning as risk-aware trajectory optimization solved as a QP and warm-started by an offline trajectory library.
More recently, many off-road systems execute learned traversability costs through sampling-based MPC, most commonly MPPI [47], to generate short-horizon actions [17, 16, 3, 41]. Such methods are flexible but can be sensitive to objective weighting and may exhibit oscillations when trading off competing terms [46].
Relation to this work: In contrast to MPPI-style short-horizon control, our method uses spatial gradients from the INR to perform gradient-based trajectory optimization that refines an entire local path and co-optimizes a traversability-aware speed profile. Compared with STEP, which warm-starts a QP using an offline trajectory library, we generate an initial path online with A and refine it online within our planning cycle.
III Proposed Method
@@
Figure 2: Our perception module takes a camera image and a LiDAR point cloud and produces an implicit representation of various terrain properties. During training, we use PyTorch’s autograd to obtain the terrain slope directly from ground max elevation $f_{\texttt{g\_max}}$, which provides an extra supervision signal. During deployment, the proposed planning and control framework incorporates gradient information provided by the implicit representation into a novel gradient-based trajectory optimization framework, where speed profile and path geometry are co-optimized.
Given an RGB image and a LiDAR point cloud from onboard sensors, the task is to infer continuous terrain attributes (e.g., elevation and slope) over a local region $x_{\min},x_{\max}]\times[y_{\min},y_{\max}]$ in the vehicle frame and use them to navigate to goals provided by some higher-level planner (which we assume to be given).
III-A Learning Implicit Representations
We train a DNN that maps images and point clouds to traversability-relevant terrain properties. A shared encoder fuses inputs into a BEV feature plane, and query-based decoders output continuous terrain attributes at arbitrary locations (Fig. 2). Unlike fixed-grid predictors [16, 3, 15], our implicit formulation uses lightweight decoders and provides spatial gradients for trajectory optimization.
III-A1 Encoder
Our encoder combines Lift-Splat-Shoot (LSS) [39] to project rich image semantics into BEV with PointPillars [25] to encode LiDAR-derived 3D structure in BEV, leveraging complementary semantic and geometric cues for more robust traversability prediction. Given an image $X\in\mathbb{R}^{C\times H\times W}$, we define d as the desired downsampling factor. We extract visual features using DINOv2 [34], obtaining $X_{\texttt{dino}}\in\mathbb{R}^{C_{\texttt{dino}}\times H_{p}\times W_{p}}$. These features are compressed with a lightweight 2D convolution block ($\operatorname{Conv}+\operatorname{GroupNorm}+\operatorname{ReLU}$) to produce $X_{\texttt{dc}}\in\mathbb{R}^{C_{\texttt{dc}}\times H_{p}\times W_{p}}$. From $X_{\texttt{dc}}$, we predict depth logits $X_{\texttt{depth}}\in\mathbb{R}^{D\times H_{d}\times W_{d}}$, where $H_{d}=H/\texttt{d}$ and $W_{d}=W/\texttt{d}$. We also interpolate $X_{\texttt{dc}}$ to obtain 2D image features $X_{\texttt{feat}}\in\mathbb{R}^{C_{\texttt{dc}}\times H_{d}\times W_{d}}$. The pair $(X_{\texttt{feat}},X_{\texttt{depth}})$ is lifted to a 3D volumetric feature of size $C_{\texttt{dc}}\times D\times H_{d}\times W_{d}$ via depth unprojection, which is then splatted onto the BEV plane, following LSS. The resulting image BEV features are compressed to $X_{\texttt{img}}\in\mathbb{R}^{C_{\texttt{img}}\times Y\times X}$. For LiDAR, we apply PointPillars to obtain point cloud BEV features $X_{\texttt{pcd}}\in\mathbb{R}^{C_{\texttt{pcd}}\times Y\times X}$. The image and point cloud BEV features are concatenated and passed through a 2D convolution block to produce $C_{\texttt{bev}}$ channels, followed by a CBAM attention module [48]. The final encoder output is a fused BEV feature map $X_{\texttt{bev}}\in\mathbb{R}^{C_{\texttt{bev}}\times Y\times X}$.
III-A2 Decoder
Let the encoder inputs be $\mathbf{X}$ (an image and a point cloud) and let the encoder output a BEV feature map $X_{\texttt{bev}}\in\mathbb{R}^{C_{\texttt{bev}}\times Y\times X}$. Given a continuous 2D query point $p=(p_{x},p_{y})$ in the local map frame, we extract a feature vector $\psi(p,\mathbf{X})\in\mathbb{R}^{C_{\texttt{bev}}}$ by bilinearly sampling $X_{\texttt{bev}}$ at $(p_{x},p_{y})$. [^1] We use differentiable bilinear interpolation (e.g., grid_sample), so $\psi(p,\mathbf{X})$ is differentiable w.r.t. $p$. Because $X_{\texttt{bev}}$ is produced by convolutional blocks with attention (CBAM), $\psi(p,\mathbf{X})$ aggregates spatial context from a neighborhood around $p$.
We represent each terrain attribute as a coordinate-conditioned MLP (an implicit/coordinate-based decoder) that takes both the local BEV feature and a positional encoding of $p$ as input [36]. Directly feeding $p$ to an MLP can underfit high-frequency structure [42]; we therefore use Fourier features (positional encoding) $\gamma(p)\in\mathbb{R}^{\texttt{ff\_dim}}$:
$\gamma(p)=\big[\sin(2\pi Bp),\ \cos(2\pi Bp)\big],$ (1)
where $B\in\mathbb{R}^{(\texttt{ff\_dim}/2)\times 2}$ is a fixed random matrix (as in [42]). This encoding improves the decoder’s ability to represent fine-scale terrain variations (e.g., small rocks or partially buried roots) that matter for traversability.
For compactness, we write the full decoding pipeline as a single function $f_{\theta}:\mathbb{R}^{2}\rightarrow\mathbb{R}$:
$f_{\theta}(p)=\mathrm{MLP}_{\theta}\big([\gamma(p);\ \psi(p,\mathbf{X})]\big),$ (2)
which is continuous and differentiable w.r.t. $p$. This allows us to query terrain attributes at arbitrary locations and to obtain spatial gradients via automatic differentiation when needed by training and trajectory optimization.
We instantiate five decoders: $f_{\texttt{g\_min}}$, $f_{\texttt{g\_max}}$, $f_{\texttt{c\_mask}}$, $f_{\texttt{c\_min}}$, $f_{\texttt{bump}}$. $f_{\texttt{g\_min}}$ and $f_{\texttt{g\_max}}$ predict minimum/maximum ground elevation; $f_{\texttt{c\_mask}}$ and $f_{\texttt{c\_min}}$ predict a ceiling mask and minimum ceiling height (used only during training, inspired by TerrainNet [32], to prevent overhanging structures from inflating ground estimates); and $f_{\texttt{bump}}$ predicts terrain bumpiness as defined in Sec. III-A3.
III-A3 Data Preparation and Training
The ground truth labels for ground elevations and ceiling predictions are generated in similar fashion to TerrainNet [32], where LiDAR scans in the time range $t-\delta t,t+\delta t]$ are aggregated using odometry to generate dense annotations at time $t$. Suppose, at a fixed time stamp, each grid has points $\{(x_{i},y_{i},z_{i})\}_{i=1}^{k}$, ordered by ascending order of $z$. The minimum ground elevation of each grid is then $\frac{1}{m}\sum^{m}_{i=1}z_{i}$, where $m$ is a parameter. An average over $m$ points is used instead of $z_{1}$ in order to smooth out noise. The minimum ceiling elevation is computed by detecting gaps in $z$: if $z_{i+1}-z_{i}>\delta z$ and $z_{j+1}-z_{j}\leq\delta z~\forall j<i$, then we classify all points $j$ with $j\geq i+1$ as part of the ceiling, and the ceiling elevation is set to $z_{i+1}$. Once the ceiling is determined, the maximum ground elevation is simply set to $z_{i}$. In cases where no ceiling is detected, the maximum ground elevation is set to $z_{k}$.
To quantify terrain bumpiness, we use the root-mean-square (RMS) of the vehicle’s vertical acceleration $a_{z}$ over a short window ($0.1\text{\,}\mathrm{s}$). Because the RMS of $a_{z}$ on the same terrain increases approximately proportionally with vehicle speed [37], we normalize the measured RMS by the vehicle speed $v$ during dataset generation. Specifically, for each time $t$ we compute $\mathrm{rms}_{z}(t)=\sqrt{\frac{1}{T}\int_{t-T}^{t}a_{z}(\tau)^{2}\,d\tau}$ (with $T=0.1$ s), form the speed-normalized value $\tilde{b}(t)=\mathrm{rms}_{z}(t)/(v(t)+\varepsilon)$, and rescale it to $(0,1)$ via $b(t)=\operatorname{sigmoid}(\tilde{b}(t))$. We use $b(t)$ as the bumpiness label (i.e. label for $f_{\texttt{bump}}$) and associate it to the corresponding map cell using odometry. In our trajectory optimization framework (Sec. III-B2), we intentionally exploit the underlying proportional relationship between vibration and speed to encourage speed adaptation based on predicted bumpiness.
To balance the bumpiness data distribution, we compute a histogram of the bumpiness values and give each bumpiness label a weight that is inversely proportional to its frequency.
Finally, for training, we leverage the differentiability of $f_{\texttt{g\_max}}$ to provide extra supervision signal using terrain slope. We use PyTorch’s autograd functionality to compute the gradient of $f_{\texttt{g\_max}}$ with respect to the query point, $\partial_{p}f_{\texttt{g\_max}}$, and compute another loss term by comparing $\partial_{p}f_{\texttt{g\_max}}$ to the ground truth terrain slope (see Figure 2). $L_{1}$ loss is used for $f_{\texttt{g\_min}},f_{\texttt{g\_max}},f_{\texttt{c\_min}},f_{\texttt{bump}}$ and $\partial_{p}f_{\texttt{g\_max}}$. Cross entropy loss is used for both ceiling mask and depth logits prediction.
III-B Planning with Implicit Representations
Our stack combines a geometric path planner with a gradient-based trajectory optimizer that uses $f_{\texttt{bump}}$ to refine the path and speed profile for smoother motion.
III-B1 Path Planner